OpenStudy (anonymous):

Solve the following system of equations by linear combination: 3x + 2y = 7 5x – 4y = 19 A. The solution is (3, –1). B. The solution is (5, 1). C. There is no solution. D. There are an infinite number of solutions.

OpenStudy (anonymous):

2(3x + 2y = 7) 5x - 4y = 19 6x + 4y = 14 5x - 4y = 19 11x = 33 x = 33/11 x = 3

OpenStudy (anonymous):

Substituting that for either equation, we have: 3(3) + 2y = 7 9 + 2y = 7 2y = -2 y = -1

OpenStudy (anonymous):

I hope this helps !